OpenStudy (anonymous):

What is the midpoint of the line segment with endpoints (5, -2) and (1, -6)? Select one of the options below as your answer: A. (2, -2) B. (3, -4) C. (3, -6) D. (4, -3)

OpenStudy (angela210793):

XM=(XA+XB)/2=(5+1)/2=3 YM=(YA+YB)/2=(-2-6)/2=-8/2=-4 M(3;-4)

OpenStudy (amistre64):

add the points together and divide by 2 :)

OpenStudy (angela210793):

\(\color{#ea93ef}{\text{T}}\)\(\color{#14e7f1}{\text{h}}\)\(\color{#34c8f7}{\text{n}}\)\(\color{#9e5ac5}{\text{x}}\)\(\color{#bca5dc}{\text{ }}\)\(\color{#187b3f}{\text{A}}\)\(\color{#c75487}{\text{m}}\)\(\color{#d756a6}{\text{i}}\)\(\color{#c77f6b}{\text{s}}\)\(\color{#3e7457}{\text{t}}\)\(\color{#ed2297}{\text{r}}\)\(\color{#3881de}{\text{e}}\)\(\color{#111884}{\text{ }}\)\(\color{#624a1c}{\text{:}}\)\(\color{#73f616}{\text{)}}\)\(\color{#ad8745}{\text{}}\)